Case 15.4: Sapphire Coffee - Part 2 This case is an extension of the Sapphire Coffee case from Chapter 14. Case the issue would be: Is the $5.00 increase in sales for each one square foot increase in store size adequate and if not would knowing the location of the store and whether it had a drive through window help explain sales. A stand alone analysis would start with the output from the case in Chapter 14 with a discussion that the low R-square value and relatively high confidence interval would justify looking for additional independent variables. This would lead to dummy variables: dl No 0 Yes 1 Near a college Drive through window No0 Yes 1 The Excel output with the dummy variables is: SUMMARY OUTPUT Regression Statistics Multiple R R Square Adjusted R Square0.71011358 0.852547827 0.726837797 Standard Error Observations 252.0293585 53 df Ss MS Regressiorn Residual Total 3 8281618.949 2760540 43.4602 7.58205E-14 49 3112421.08 63518.8 52 11394040.03 Coefficients Standard Error titat P-value Lower 95% Upper 95% 1481.063116 273.490089 5.342289 2.37E-06 911.4645391 2010.662 Store Size (Sq. Ft.) 3.256333342 0.289857069 11.23427 3.68E-15 2.673844106 3.838823 -109.3628452 84.63282732 1.2922 0.202348 -279.4387761 60.71309 314 3532146 105.6491529 2.975445 0.004533 102.0434219 526.663 Intercept College Nearby Drive Thru Adding the dummy variables increased the R-square value of the model, but knowing if there is a college near-by is not significant. Having a drive through window seems to significantly increase sales but how much is somewhat questionable given the wide confidence interval.
Multiple Regression #2 In the orking for a local retail store you have developed ing estimated regression equation: y= 22,16 7 412 +8182-93x3-71x eekly sales Xi = Local unemployment rate kly average high temperature Number of activities in the local community Average gasoline price a. Interpret the values of b. b2. b3. and b, in this ted regression equation. b. What is the estimated sales if the unemployment rate is 5.7%, the average high temperature is 61°, there are 14 activities, and gasoline average price
Answer #1

15.4) As per the excel results given above having the dummy variable drive through window,

Increase in one drive through window increases the sales by $314.

Problem 2:-

given equation y = 22,167 - 412x1 + 818x2 - 93x3 - 71x4

this equation can be related to

y = b0 +b1x1+b2x2+b3x3+b4x4

a) as per the given data above the coefficients of variables can be explained as below

b1 = With increase in 1 unit of local unemployment rate (x1) weekly sales (y) decreases by 412 units.

b2 = With increase in 1 unit of weekly average high temperature (x2) weekly sales (y) increases by 818 units.

b3 = With increase in 1 unit of number of activities in the local community (x3) weekly sales (y) decreases by 93 units.

b4 = With increase in 1 unit of Average gasoline price (x4) weekly sales (y) decreases by 71 units.

b) given x1 =5.7% x2 = 61o x3 =14 x4 = $1.39

y = 22,167 - 412(5.7) + 818 (61) - 93 (14) - 71(1.39) = 68315.91 units